# Gradient flows of $(K,N)$-convex functions with negative $N$

### v1

@@ 1  Introduction
The theory of gradient flows was developed to generalize the classical notion of gradient-driven evolution beyond smooth and linear spaces, such as Hilbert spaces, to broader settings like Banach or metric spaces. One of the most widely studied and influential theories, with significant applications to various types of PDEs, involves the evolution of systems within a Hilbert space $X$, driven by a lower semicontinuous and convex (or $K$-convex) functional $\varphi\colon X\to(-\infty,+\infty]$ with non-empty, proper domain $\mathcal{D}(\varphi)\coloneqq\{y\in X:\varphi(y)<\infty\}$. In this case, the evolution can be described by a locally Lipschitz curve $u\colon(0,+\infty)\to\mathcal{D}(\varphi)$ solving the differential inclusion
$u^{\prime}(t)\in-\partial\varphi\big{(}u(t)\big{)}\quad\text{ for }\mathcal{L}^{1}\text{-a.e. }t>0\quad\text{ with }\quad\lim_{t\downarrow 0}u(t)=u_{0}\in\overline{\mathcal{D}(\varphi)},$ (1)
@@
$\frac{\mathsf{d}}{\mathsf{d}t}\varphi\big{(}u(t)\big{)}=-\left\lVert u^{\prime}(t)\right\rVert_{X}^{2}=-\left\lVert\partial\varphi\big{(}u(t)\big{)}\right\rVert^{2}_{X},\qquad\text{ for }\mathcal{L}^{1}\text{-a.e.}t>0,$ (2)
@@
For any sufficiently small time step $\tau>0$, (3) recursively defines a sequence $\left(U^{n}_{\tau}\right)_{n\in\mathbb{N}}$, which in turns induces a piecewise constant interpolant
$\bar{U}_{\tau}(t)\coloneqq U^{n}_{\tau}=J^{n}_{\tau}(u_{0})\text{ if }t\in\big{(}(n-1)\tau,n\tau\big{]},$ (4)
@@
$|D^{-}\varphi|(u)\coloneqq\limsup_{z\to y}\dfrac{\big{(}f(z)-f(y)\big{)}^{-}}{\mathsf{d}(z,y)},$
@@
$\dfrac{\mathsf{d}}{\mathsf{d}t}\varphi\big{(}u(t)\big{)}=|\dot{u}|^{2}(t)=-|D^{-}\varphi|^{2}\big{(}u(t)\big{)}\text{ for }\mathcal{L}^{1}\text{-a.e. }t>0,$
@@
$\dfrac{1}{2}\dfrac{\mathsf{d}}{\mathsf{d}t}\left\lVert u(t)-w\right\rVert_{X}^{2}=\langle u^{\prime}(t),u(t)-w\rangle_{X}$
@@
$\dfrac{1}{2}\dfrac{\mathsf{d}}{\mathsf{d}t}\left\lVert u(t)-w\right\rVert_{X}^{2}\leq\varphi(w)-\varphi(u(t))-\dfrac{K}{2}\left\lVert u(t)-w\right\rVert_{X}^{2}\quad\text{ for every }w\in\mathcal{D}(\varphi).$
@@
$\dfrac{1}{2}\dfrac{\mathsf{d}}{\mathsf{d}t}\mathsf{d}^{2}(u(t),w)+\dfrac{K}{2}\mathsf{d}^{2}(u(t),w)\leq\varphi(w)-\varphi(u(t))\quad\text{ for every }w\in\mathcal{D}(\varphi).$
@@
The study of EVI gradient flows has played a central role in characterizing the synthetic curvature-dimension condition in the context of metric measure spaces. In particular, it has been essential in linking the Eulerian and Lagrangian approaches within this theory. A first step in this direction was taken in [AGS14b], where it is shown that, in a $\mathsf{RCD}(K,\infty)$ space, the gradient flow of the relative entropy exists for suitable initial data and satisfies the $\operatorname{EVI}_{K}$ condition.
@@
The main goal of this paper is to study the fundamental properties of $\operatorname{EVI}_{K,N}$ gradient curves for negative values of the parameter $N$, associated with lower semicontinuous functionals defined on a metric space. In particular, in Section 4 we prove the uniqueness (assuming existence) of $\operatorname{EVI}_{K,N}$ gradient curves starting from arbitrary points in the domain, while in Section 5, we carefully examine the relationship between the existence of $\operatorname{EVI}_{K,N}$ gradient curves and the existence of curves of maximal slope that satisfy either the Energy Dissipation Equality (EDE) or the Energy Dissipation Inequality (EDI). The framework we adopt extends that of [Oht16] by defining $(K,N)$-convexity in a manner that includes functions taking the value $\infty$- ∞. This generalization is motivated by specific examples of $(K,N)$-convex functions defined on open intervals of $\mathbb{R}$, whose natural extensions to the boundary points take the value $\infty$- ∞. For instance, for any $N<0$ the function $f\colon x\mapsto-N\log x$ is $(0,N)$-convex in $(0,+\infty)$ and $\lim_{x\to 0}f(x)=-\infty$. This will be a crucial point when studying the $\operatorname{EVI}_{K,N}$-gradient flow of the Boltzmann entropy in ${\mathsf{C}D}^{e}(K,N)$-spaces, namely metric measure spaces satisfying the entropic curvature-dimension condition. Furthermore this example illustrates that we cannot generally expect a function $\varphi$, $(K,N)$-convex for some $N<0$, to be coercive in the sense of [AGS08, Property 2.1b], meaning that for each $\tau>0$ and $x\in X$ it holds
@@
We conclude by recalling that the range of values for the dimension $N$ has been extended to include negative values also in the theory of $\mathsf{CD}(K,N)$ spaces, as presented in [Oht16]. This extension was motivated in the setting of weighted Riemannian manifolds by the work [OT11], where the convexity of a certain generalization of the relative entropy (inspired by information theory) was characterized combining the condition $\text{Ric}_{N}\geq 0$ with the convexity of another weight function, with $N$ possibly taking negative values depending on the choice of entropy. In fact, most of the results on this topic are obtained in the case of weighted Riemannian manifolds. For example, in the Euclidean setting a direct application of the results in [BL76] ensures that given any convex measure $\mu$ with full dimensional convex support and $C^{2}$ density $\Psi$, the space $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl.}},\mu)$ satisfies the $\mathsf{CD}(0,N)$ condition for $1/N\in(-\infty,1/n]$ (i.e. $N\in(-\infty,0)\cup[n,\infty]$), in the sense that $\mathrm{Ric}_{N}\geq 0$. This class of measures, introduced by Borell in [Bor75], extends the set of the so-called log-concave measures and it has been largely studied for example in [Bob07, BL09, Kol14]. In particular, following the terminology adopted by Bobkov, the case $N\in(-\infty,0)$ corresponds to the “heavy-tailed measures” (see also [BCR05]), identified by the condition that $1/\Psi^{1/(n-N)}$ is convex. An explicit example of these measures is given by the family of Cauchy probability measures on $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl.}})$
$\mu^{n,\alpha}=\dfrac{c_{n,\alpha}}{\big{(}1+\left\lvert x\right\rvert^{2}\big{)}^{\frac{n+\alpha}{2}}}\,\mathsf{d}x,\quad\alpha>0,$ (6)
where $c_{n,\alpha}>0$ is a normalization constant. It then follows that $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl.}},\mu^{n,\alpha})$ is a $\mathsf{CD}(0,-\alpha)$ space. Moreover, in [Mil17] it is proved that the $n$-dimensional unit sphere $\mathbb{S}^{n}$ equipped with the harmonic measure, namely the hitting distribution by the Brownian motion started at $x\in\mathbb{S}^{n}$, (which can be equivalently described as the probability measure whose density is proportional to $\mathbb{S}^{n}\ni y\to 1/|y-x|^{n+1}$) is a $\mathsf{CD}\big{(}n-1-(n+1)/4,-1\big{)}$ space. More generally, Milman provides an analogue to the family (6) of Cauchy measures in $\mathbb{R}^{n+1}$, showing that the family of probability measures on $\mathbb{S}^{n}$ having density proportional to
@@
Following [Oht16], the study of nonsmooth $\mathsf{CD}(K,N)$ spaces has continued in [MRS23, MR23], where several important properties have been established. In particular, these works identify the class of metric measure spaces equipped with quasi-Radon measure as the natural framework for defining the $\mathsf{CD}(K,N)$ condition. Drawing an analogy with the positive-dimensional case $N>0$, our study of $\operatorname{EVI}_{K,N}$ gradient flow may provide a fundamental tool for bridging the Eulerian and Lagrangian approaches on lower curvature bounds. This, in turn, could yield a consistent definition of $\mathsf{RCD}(K,N)$ spaces for negative values of $N$, allowing the development of an analytic approach to the study of metric measure spaces equipped with singular (quasi-Radon) measures.
@@ 2. $(ii)$
$f(\gamma_{t})\leq(1-t)f(x_{0})+tf(x_{1})-\tfrac{\lambda}{2}t(1-t)\,\mathsf{d}(x_{0},x_{1})^{2}\qquad t\in[0,1].$ (8)
@@
$\operatorname{Hess}f(v,v)\geq K|v|^{2}\qquad\text{for all }v\in TM.$ (9)
@@ 2.1  $(K,N)$-convex functions for $N<0$
These distortion coefficients have the following monotonicity properties: for every fixed $t\in[0,1]$ it holds
@@ 4. $(d)$
    for $K<0$, the function $x\mapsto-N\log\big{(}\cos(x\sqrt{K/N})\big{)}$ on $\big{(}-\tfrac{\pi}{2}\sqrt{N/K},\tfrac{\pi}{2}\sqrt{N/K}\big{)}$.
@@
$f_{N}(c)\leq\frac{\sigma_{K,N}^{(\frac{c-b}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}{1-\sigma_{K,N}^{(\frac{b-a}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}f(a)+\frac{\sigma_{K,N}^{(\frac{c-b}{d-b})}(d-b)}{1-\sigma_{K,N}^{(\frac{b-a}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}f(b).$ (18)
@@
The same identity holds by explicit computation also for $K=0$, and, as a consequence, we deduce that
@@ 2.2  Gradient flows
In the following we will usually denote by $(y_{t})_{t\in(0,T)}$ the curves in $X$ defined on $(0,T)$. Moreover, we say that a curve $(y_{t})_{t\in(0,T)}$ starts at $\bar{y}\in X$ if $\lim_{t\to 0}y_{t}=\bar{y}$.
@@
The following result highlights an important property of $\operatorname{EVI}_{\lambda}$ gradient curves, namely the so-called $\lambda$-contractivity. As an immediate consequence, this implies the uniqueness of $\operatorname{EVI}_{\lambda}$ gradient curves for a given initial condition. For a detailed proof of this result, we refer to [MS20, Thm. 3.5]. It is worth noting that in our paper, the definition of $\operatorname{EVI}_{\lambda}$ gradient curves slightly differs from the one in [MS20], but the same proof arguments apply in our context as well.
@@
Let $g\colon X\to\bar{\mathbb{R}}$ be a lower semicontinuous functional and $\lambda\in\mathbb{R}$. Let $(z^{1}_{t})_{t\in(0,T)},(z^{2}_{t})_{t\in(0,T)}\subset\mathcal{D}[g]$ be two $\operatorname{EVI}_{\lambda}$ gradient curves for $g$. Then for every $0<r<s<T$ it holds
@@
In particular, for every $z_{0}\in\mathcal{D}[g]$ there exists at most one $\operatorname{EVI}_{\lambda}$ gradient curves for $g$ satisfying the initial condition $\lim_{t\downarrow 0}z_{t}=z_{0}$.
@@
ensuring in particular that $z^{\prime}\in\mathcal{D}[g]$. Observe that the triangular inequality guarantees that $\mathsf{d}(z_{s},z)\leq\mathsf{d}(z_{s},z^{\prime})+\mathsf{d}(z^{\prime},z)$ for every $s\in(0,T)$, therefore we deduce that
$\begin{split}\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\frac{\mathsf{d}(z_{t},z)^{2}}{2}=\mathsf{d}(z_{t},z)\,\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\mathsf{d}(z_{t},z)&\leq\mathsf{d}(z_{t},z)\,\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\mathsf{d}(z_{t},z^{\prime})=\frac{1}{r}\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\frac{\mathsf{d}(z_{t},z^{\prime})^{2}}{2}.\end{split}$ (33)

### v2

# cell-space-limits = 1pt

@@ 1  Introduction
The theory of gradient flows was developed to generalize the classical notion of gradient-driven evolution beyond smooth and linear spaces, such as Hilbert spaces, to broader settings like Banach or metric spaces. One of the most widely studied and influential theories, with significant applications to various types of PDEs, involves the evolution of systems within a Hilbert space $X$, driven by a lower semicontinuous and convex (or $K$-convex) functional $\varphi\colon X\to(-\infty,+\infty]$ with non-empty, proper domain $\mathcal{D}[\varphi]\coloneqq\{y\in X:\varphi(y)<\infty\}$. In this case, the evolution can be described by a locally Lipschitz curve $u\colon(0,+\infty)\to\mathcal{D}[\varphi]$ solving the differential inclusion
$u^{\prime}(t)\in-\partial\varphi\big{(}u(t)\big{)}\quad\text{ for }\mathcal{L}^{1}\text{-a.e. }t>0\quad\text{ with }\quad\lim_{t\downarrow 0}u(t)=u_{0}\in\overline{\mathcal{D}[\varphi]},$ (1)
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\varphi\big{(}u(t)\big{)}=-\left\lVert u^{\prime}(t)\right\rVert_{X}^{2}=-\left\lVert\partial\varphi\big{(}u(t)\big{)}\right\rVert^{2}_{X},\qquad\text{ for }\mathcal{L}^{1}\text{-a.e. }t>0,$ (2)
@@
For any sufficiently small time step $\tau>0$, (3) recursively defines a sequence $\{U^{n}_{\tau}\}_{n\in\mathbb{N}}$, which in turns induces a piecewise constant interpolant
$\bar{U}_{\tau}(t):=U^{n}_{\tau}=J^{n}_{\tau}(u_{0})\text{ if }t\in\big{(}(n-1)\tau,n\tau\big{]},$ (4)
@@
$|D^{-}\varphi|(u):=\limsup_{z\to y}\dfrac{\big{(}f(z)-f(y)\big{)}^{-}}{\mathsf{d}(z,y)},$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\varphi\big{(}u(t)\big{)}=|\dot{u}|^{2}(t)=-|D^{-}\varphi|^{2}\big{(}u(t)\big{)}\text{ for }\mathcal{L}^{1}\text{-a.e. }t>0,$
@@
$\dfrac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\left\lVert u(t)-w\right\rVert_{X}^{2}=\langle u^{\prime}(t),u(t)-w\rangle_{X}$
@@
$\dfrac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\left\lVert u(t)-w\right\rVert_{X}^{2}\leq\varphi(w)-\varphi(u(t))-\dfrac{K}{2}\left\lVert u(t)-w\right\rVert_{X}^{2}\quad\text{ for every }w\in\mathcal{D}(\varphi).$
@@
$\dfrac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{d}^{2}(u(t),w)+\dfrac{K}{2}\mathsf{d}^{2}(u(t),w)\leq\varphi(w)-\varphi(u(t))\quad\text{ for every }w\in\mathcal{D}(\varphi).$
@@
The study of EVI gradient flows has played a central role in characterizing the synthetic curvature-dimension condition in the context of metric measure spaces. In particular, it has been essential in linking the Eulerian and Lagrangian approaches within this theory. A first step in this direction was taken in [GKO13], where the correct link between the ${\sf CD}(K,\infty)$ condition and the Bochner inequality was established in the setting of Alexandrov spaces. Building on the ideas introduced in that work, [AGS14b] subsequently showed that, in an $\mathsf{RCD}(K,\infty)$ space, the gradient flow of the relative entropy exists for suitable initial data and satisfies the $\operatorname{EVI}_{K}$ condition. In turn, this result allows one to deduce the Bakry–Émery formulation of the $\mathsf{RCD}(K,\infty)$ condition, and ultimately to derive the Bochner inequality in a weak sense.
@@
The main goal of this paper is to study the fundamental properties of $\operatorname{EVI}_{K,N}$ gradient curves for negative values of the parameter $N$, associated with lower semicontinuous functionals defined on a metric space. In particular, in Section 4 we prove the uniqueness (assuming existence) of $\operatorname{EVI}_{K,N}$ gradient curves starting from arbitrary points in the domain, while in Section 5, we carefully examine the relationship between the existence of $\operatorname{EVI}_{K,N}$ gradient curves and the existence of curves of maximal slope that satisfy either the Energy Dissipation Equality (EDE) or the Energy Dissipation Inequality (EDI). The framework we adopt extends that of [Oht16] by defining $(K,N)$-convexity in a manner that includes functions taking the value $\infty$- ∞. This generalization is motivated by specific examples of $(K,N)$-convex functions defined on open intervals of $\mathbb{R}$, whose natural extensions to the boundary points take the value $\infty$- ∞. For instance, for any $N<0$ the function $f\colon x\mapsto-N\log x$ is $(0,N)$-convex in $(0,+\infty)$ and $\lim_{x\to 0}f(x)=-\infty$. In this case, the gradient flow of $f$ satisfies $\dot{y}_{t}=-f^{\prime}(y_{t})=-N/y_{t}$. Thus, for every starting point $y_{0}\geq 0$, we have $y_{t}=\sqrt{y_{0}^{2}+2Nt}$ for $t\in\big{[}0,-\tfrac{y_{0}^{2}}{2N}\big{)}$ and $y_{t}=0$ for $t>-\tfrac{y_{0}^{2}}{2N}$.
This will be a crucial point when studying the $\operatorname{EVI}_{K,N}$-gradient flow of the Boltzmann entropy in ${\mathsf{C}D}^{e}(K,N)$-spaces, namely metric measure spaces satisfying the entropic curvature-dimension condition. Furthermore this example illustrates that we cannot generally expect a function $\varphi$, $(K,N)$-convex for some $N<0$, to be coercive in the sense of [AGS08, Property 2.1b], meaning that for each $\tau>0$ and $x\in X$ it holds
@@
We conclude by recalling that the range of values for the dimension $N$ has been extended to include negative values also in the theory of $\mathsf{CD}(K,N)$ spaces, as presented in [Oht16]. This extension was motivated in the setting of weighted Riemannian manifolds by the work [OT11, OT13], where the convexity of a certain generalization of the relative entropy (inspired by information geometry, see [OT11, §3.1]) was characterized combining the condition $\text{Ric}_{N}\geq 0$ with the convexity of another weight function, with $N$ possibly taking negative values depending on the choice of entropy. Indeed, the entropy functionals in [OT11] belong to the class of Bregman divergences, which are dissimilarity functionals encoding dually flat Riemannian structures on spaces of probability densities, in the sense of information geometry (see e.g. [Ama16, §1.3.2, and p. 28]).
In fact, most of the results on this topic are obtained in the case of weighted Riemannian manifolds. For example, in the Euclidean setting, a direct application of the results in [BL76] ensures that given any convex measure $\mu$ with full dimensional convex support and $C^{2}$ density $\Psi$, the space $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl}},\mu)$ satisfies the $\mathsf{CD}(0,N)$ condition for $1/N\in(-\infty,1/n]$ (i.e., $N\in(-\infty,0)\cup[n,\infty]$), in the sense that $\mathrm{Ric}_{N}\geq 0$. This class of measures, introduced by Borell in [Bor75], extends the set of the so-called log-concave measures and it has been largely studied for example in [Bob07, BL09, Kol14]. In particular, following the terminology adopted by Bobkov, the case $N\in(-\infty,0)$ corresponds to the “heavy-tailed measures” (see also [BCR05]), identified by the condition that $1/\Psi^{1/(n-N)}$ is convex. An explicit example of these measures is given by the family of Cauchy probability measures on $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl}})$
$\mu^{n,\alpha}=\dfrac{c_{n,\alpha}}{\big{(}1+\left\lvert x\right\rvert^{2}\big{)}^{\frac{n+\alpha}{2}}}\,\,\mathrm{d}x,\quad\alpha>0,$ (6)
where $c_{n,\alpha}>0$ is a normalization constant. It then follows that $(\mathbb{R}^{n},\mathsf{d}_{\text{Eucl}},\mu^{n,\alpha})$ is a $\mathsf{CD}(0,-\alpha)$ space. Moreover, in [Mil17] it is proved that the $n$-dimensional unit sphere $\mathbb{S}^{n}$ equipped with the harmonic measure, namely the hitting distribution by the Brownian motion started at $x\in\mathbb{S}^{n}$, (which can be equivalently described as the probability measure whose density is proportional to $\mathbb{S}^{n}\ni y\to 1/|y-x|^{n+1}$) is a $\mathsf{CD}\big{(}n-1-(n+1)/4,-1\big{)}$ space. More generally, Milman provides an analogue to the family (6) of Cauchy measures in $\mathbb{R}^{n+1}$, showing that the family of probability measures on $\mathbb{S}^{n}$ having density proportional to
@@
Following [Oht16], the study of non-smooth $\mathsf{CD}(K,N)$ spaces has continued in [MRS23, MR23], where several important properties have been established. In particular, these works identify the class of metric measure spaces equipped with a quasi-Radon measure as the natural framework to define the $\mathsf{CD}(K,N)$ condition. Drawing an analogy with the positive-dimensional case $N>0$, our study of $\operatorname{EVI}_{K,N}$ gradient flow may provide a fundamental tool for bridging the Eulerian and Lagrangian approaches on lower curvature bounds. This, in turn, could yield a consistent definition of $\mathsf{RCD}(K,N)$ spaces for negative values of $N$, allowing the development of an analytic approach to the study of metric measure spaces equipped with singular (quasi-Radon) measures.
@@ 2. $(ii)$
$f(\gamma_{t})\leq(1-t)f(x_{0})+tf(x_{1})-\frac{\lambda}{2}t(1-t)\,\mathsf{d}(x_{0},x_{1})^{2}\qquad t\in[0,1].$ (8)
@@
$\operatorname{Hess}f(v,v)\geq\lambda|v|^{2}\qquad\text{for all }v\in TM.$ (9)
@@ 2.1  $(K,N)$-convex functions for $N<0$
These distortion coefficients have the following monotonicity properties (see [Stu06]): for every fixed $t\in[0,1]$ it holds
@@ 4. $(d)$
    for $K<0$, the function $x\mapsto-N\log\big{(}\cos(x\sqrt{K/N})\big{)}$ on $\big{(}-\frac{\pi}{2}\sqrt{N/K},\frac{\pi}{2}\sqrt{N/K}\big{)}$.
@@
$f_{N}(c)\leq\frac{\sigma_{K,N}^{(\frac{c-b}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}{1-\sigma_{K,N}^{(\frac{b-a}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}f_{N}(a)+\frac{\sigma_{K,N}^{(\frac{c-b}{d-b})}(d-b)}{1-\sigma_{K,N}^{(\frac{b-a}{c-a})}(c-a)\,\sigma_{K,N}^{(\frac{d-c}{d-b})}(d-b)}f_{N}(b).$ (18)
@@
The same identity holds by explicit computation also for $K=0$, indeed
$(c-a)(d-b)-(b-a)(d-c)=(d-a)(c-b).$ (20)
As a consequence, we deduce that
@@ 2.2  Gradient flows
In the following we will usually denote by $(y_{t})_{t\in(0,T)}$ the curves in $X$ defined on $(0,T)$ and, given a function $f\colon I\to\mathbb{R}$ defined on some interval $I$, we denote by $\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}f(t)$ the right derivative of $f$ at the point $t\in I$, namely
$\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}f(t)\coloneqq\limsup_{h\downarrow 0}\dfrac{f(t+h)-f(t)}{h}.$
Moreover, we say that a curve $(y_{t})_{t\in(0,T)}$ starts at $\bar{y}\in X$ if $\lim_{t\downarrow 0}y_{t}=\bar{y}$.
@@
The following result highlights an important property of $\operatorname{EVI}_{\lambda}$ gradient curves, namely the so-called $\lambda$-contractivity. As an immediate consequence, this implies the uniqueness of $\operatorname{EVI}_{\lambda}$ gradient curves for a given initial condition.
@@
Let $g\colon X\to\bar{\mathbb{R}}$ be a proper lower semicontinuous functional and $\lambda\in\mathbb{R}$. Let $(z^{1}_{t})_{t\in(0,T)},(z^{2}_{t})_{t\in(0,T)}\subset\mathcal{D}[g]$ be two $\operatorname{EVI}_{\lambda}$ gradient curves for $g$. Then for every $0<r<s<T$ it holds
@@
In particular, for every $z_{0}\in\overline{\mathcal{D}[g]}$ there exists at most one $\operatorname{EVI}_{\lambda}$ gradient curve $(z_{t})_{t\in(0,T)}$ for $g$ satisfying the initial condition $\lim_{t\downarrow 0}z_{t}=z_{0}$.
Proof.
See [MS20, Thm. 3.5]. ∎
@@
ensuring in particular that $z^{\prime}\in\mathcal{D}[g]$. Observe that, since $z^{\prime}$ lies on a geodesic connecting $z_{t}$ and $z$, we have that $\mathsf{d}(z_{t},z)=\mathsf{d}(z_{t},z^{\prime})+\mathsf{d}(z^{\prime},z)$. Moreover, the triangular inequality guarantees that $\mathsf{d}(z_{s},z)\leq\mathsf{d}(z_{s},z^{\prime})+\mathsf{d}(z^{\prime},z)$ for every $s\in(0,T)$, therefore we deduce that
$\begin{split}\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\frac{\mathsf{d}(z_{t},z)^{2}}{2}=\mathsf{d}(z_{t},z)\,\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\mathsf{d}(z_{t},z)&\leq\mathsf{d}(z_{t},z)\,\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\mathsf{d}(z_{t},z^{\prime})=\dfrac{\mathsf{d}(z_{t},z^{\prime})}{r}\,\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\mathsf{d}(z_{t},z^{\prime})=\frac{1}{r}\frac{\,\mathrm{d}^{+}}{\,\mathrm{d}t}\frac{\mathsf{d}(z_{t},z^{\prime})^{2}}{2}.\end{split}$ (34)